OpenStudy (lucyluv16):

Hi I really don't understand, sigma notation can someone help explain it to me please? @unkleRhaukus, @agent0smith @Photon336??

OpenStudy (welshfella):



OpenStudy (lucyluv16):

\[\sum_{n=1}^{14}3n+2\]

OpenStudy (welshfella):

SIGMA NOTATION:_ Sigma is the capital Greek letter used for 'SUM' the above example means the 'Sum of the squares of the 4 numbers from 1 to 4' = 1^2 + 2^2 + 3^2 + 4^2

OpenStudy (lucyluv16):

okay I understand that but the problem i'm having is solving one of the problems in the practice worksheet i'm doing, I posted the problem up above. @welshfella

OpenStudy (lucyluv16):

please I really don't understand were i'm messing up at

OpenStudy (welshfella):

when n = 1, 3n + 2 = 3(1) + 2 = 5 the values for 2-14 can be calculated in the same way.

OpenStudy (lucyluv16):

okay but what about the 14 over sigma thing, how would I solve that????

OpenStudy (welshfella):

the 2 values around the sigma merely mean the the value on n can be 1 up to 14 and the ssigma indicates that you have to add up the values of 3n + 2 whree n = 1 up to 14.

OpenStudy (welshfella):

you can work it out this way total = 3(1 + 2 +3 +...........14) + 14(2)

OpenStudy (lucyluv16):

ok so I would add everything in the parenthesis, times it by 3 then add it to 14*2???

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

this noitation is an example of an 'Arithmetic Series' where each term is incremented by a constant value called the common difference. the first few terms are 5 , 8, 11, 14 - note the common difference is 3

OpenStudy (lucyluv16):

so would the answer would be 343??

OpenStudy (welshfella):

yes

OpenStudy (lucyluv16):

awesome! can you help me on a few more so I know i'm understanding this correctly??? please!

OpenStudy (welshfella):

OK but plz make new posts.